OpenStudy (butterflyhope):

Urgent *will medal*: The number of pizza meals sold by the concession stand was 3 times the number of hot dog meals. If the pizza meal cost $4, and the hot dog meal cost $2, and the total revenue from the meals was $658, how many of each meal were sold? Thanks c:

OpenStudy (igreen):

Hmm..I think they want you to make a system of equations. 2x + 4y = 658 x = 3y Where x = pizza meals, and y = hot dog meals.

OpenStudy (igreen):

Sorry that's: 4x + 2y = 658 x = 3y

OpenStudy (igreen):

Graphing it will give us our solution. https://www.desmos.com/calculator/j4u3ujyk7h What is the solution? @ButterflyHope

OpenStudy (igreen):

The solution is the point where the two lines intersect.

OpenStudy (igreen):

Check: Well, the solution is (141, 47) That means there were 141 pizza meals and 47 hot dog meals. Check: Pizza meals are 3 times the number of hot dog meals. Hot dog meals = 47 47 * 3 = 141. Yep, it equals 141. Pizza meals cost $4 Hot Dog meals cost $2 Total cost was 658. 4 * 141 = 564 2 * 47 = 94 564 + 94 = 658 Yep, it equals 658.

OpenStudy (butterflyhope):

Thanks :)